OpenStudy (anonymous):

take the derivative of g(x)=(3x-2)^10 * (5x^2 - x +1)^12

OpenStudy (anonymous):

\[\frac{ d }{ dx } (3x-2)^{10}*(5x^{5}-x+1)^{12}\] \[=10(3x-2)^{9}*3*(5x^{5}-x+1)^{12} +\] \[ (3x-2)^{10}* 12*(5x^{5}-x+1)^{11}*(25x^{4}-1)\]